I can't believe I don't remember this...old age LOL I thought I had the answer but I guess not. The equation is:
123 - x = 118
1+Y = 16
X:Y = ?

Answer by Maths68(1474)

123-x=118
123-x-123=118-123
Cancel out +123 and - 123
-x=5
Multiply by (-1)
-x*-1=5*-1
x=-5


1+y=16
1+y-1=16-1
Cancel out +1 and - 1
y=15

x:y = 5:15
x:y = 5/15
x/y=1/3
x:y=1:3